OpenStudy (anonymous):

solve for x: sin ((pi/4)x^2) = 1

myininaya (myininaya):

first solve: sin(u)=1 u=pi/2 but we could keep going a round and a round the circle so u=pi/2+2npi n is an integer so if sin(u)=1 when u=pi/2+2npi then sin((pi/4)x^2)=1 when (pi/4)x^2=pi/2+2npi

myininaya (myininaya):

solve that last equation for x and woolah

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

im confused! how do i solve that last eqn?

myininaya (myininaya):

using algebra multiply by 4/pi on both sides then take the square root of both sides don't forget plus or minus

OpenStudy (anonymous):

is "woolah" arkansian for voila?

myininaya (myininaya):

lol

myininaya (myininaya):

its my new word

OpenStudy (anonymous):

too much barbecue and beer for me.

OpenStudy (anonymous):

i will use it next week

myininaya (myininaya):

you can use it i guess

OpenStudy (anonymous):

lol whats the answer? :P

OpenStudy (anonymous):

i will say it like inspector clouseau

OpenStudy (anonymous):

(sqrt) 2 + 8n ???